How many atoms are in a sample containing four moles of carbon
1
Expert's answer
2020-12-15T07:53:29-0500

"1mol of C=6.023\times 10^{23} atoms"

"4moles=?"

"=4\times 6.023\times 10^{23}"

"2.469\times 10^{24}" Atoms of Carbon